Church Inc. is presently enjoying relatively high growth because of a surge in the demand for its new product. Management expects earnings and dividends to grow at a rate of 16% for the next 4 years, after which competition will probably reduce the growth rate in earnings and dividends to zero, i.e., g = 0. The company's last dividend, D, was $1.25, its beta is 1.20, the market risk premium is 5.50%, and the risk-free rate is 3.00%. What is the current price of the common stock?
a. $18.35
b. $21.01
c. $22.11
d. $16.81
e. $17.03
1
Expert's answer
2012-11-29T08:51:52-0500
Solution:
Today's price equal:
P = D*(1-g)/(k-g) = D*(1-g)/(rfr+beta(rm - rfr)-g)= $1.25*1/(0.03+1.2(0.055))=
$1.25/0.096 = $13.02

Answer: $13.02